OpenStudy (anonymous):

ILL FAN & GIVE YOU MEDAL!!!!!!! It takes Brian 15 hours longer to build a model car than it takes John. If they work together, they can build the model car in 4 hours. explain each step in figuring out how to determine the time it would take Brian to build the car on his own.

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

and here comes mathbreaker :/

OpenStudy (anonymous):

if it takes john x hours, it takes brian x + 15. hence john does 1/x of the work every hour, but brian takes 1/(x+15) Oh sorry saber, ill give you this one haha ;)

OpenStudy (anonymous):

:/

OpenStudy (anonymous):

You want to take this? :| or should i :P

OpenStudy (anonymous):

go on

OpenStudy (anonymous):

u kinda confused me with ure answer are u sure its right?

OpenStudy (johnweldon1993):

He's correct so far @xrosaber

OpenStudy (anonymous):

idk im not tht good in maths

OpenStudy (anonymous):

when it comes to word problems i suk

OpenStudy (anonymous):

so i set up the equation 1/x + 1/x+5 = 1/4?

OpenStudy (anonymous):

Yes thats right bubbles :)

OpenStudy (anonymous):

Solve for x, and just add 15 to get brians time :)

OpenStudy (johnweldon1993):

Right...because John = 1 car in 'x' hours Brian = 1 car in 'x + 15' hours... so together ...if it takes 4 hours we have \[\large \frac{1}{x} + \frac{1}{x + 15} = \frac{1}{4}\]

OpenStudy (anonymous):

makes more sense i neeed to improve in word problems :(

OpenStudy (anonymous):

oh i thought you would put 1/15x to get brains time o: & Next i find the LCM correct?

OpenStudy (anonymous):

yes and youll get 8x + 60 = x^2 + 15x x^2 + 7x + 60 = 0... remember that when you take quadratic, take the positive answer always cause time cannot be negative

OpenStudy (johnweldon1993):

Check you equation again @Mathbreaker

OpenStudy (anonymous):

i thought the LCM would be 4x(x+5)......

OpenStudy (johnweldon1993):

\[\large 8x + 60 = x^2 + 15x\] becomes \[\large x^2 + 7x - 60 = 0\]

OpenStudy (anonymous):

im so confused

OpenStudy (johnweldon1993):

@bubbles-are-cool. That is the LCM :) 4x(x + 15) remember it is 15 not 5 :)

OpenStudy (anonymous):

Haha sorry i did that mentally so i missed that :P Im sorry, its x^2 + 7x -60 = 0

OpenStudy (johnweldon1993):

So we have \[\large \frac{1}{x} + \frac{1}{x + 15} = \frac{1}{4}\] Find the LCM (Least common denominator) which as @bubbles-are-cool. has said it is 4x(x + 15) What do we do next?

OpenStudy (anonymous):

guyz wht grade math is this

OpenStudy (anonymous):

Algebra 2 x: & next youd get the common denominator ?

OpenStudy (anonymous):

This would be grade 7-9 math :)

OpenStudy (anonymous):

-.-

OpenStudy (anonymous):

honestly

OpenStudy (anonymous):

im in grade 10 havent even been close to studying stuff like that

OpenStudy (anonymous):

im in 10th too

OpenStudy (anonymous):

which school

OpenStudy (anonymous):

flvs

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

Wow... im in grade 10 you should come to india we have already started differentiation and integration. india studies are fun. but they're only for people who like it :P

OpenStudy (anonymous):

i guess we have different curricualm

OpenStudy (anonymous):

im horrible at math x.x

OpenStudy (johnweldon1993):

Right...since that is 4x(x + 15)....multiply that to the 2 fractions you have on the left...and dont forget to multiply it to the other side too... \[\large \frac{4x(x + 15)}{x} + \frac{4x(x + 15)}{x + 15} = \frac{4x(x + 15)}{4}\] becomes \[\large 4(x + 15) + 4x = x(x + 15)\] Now what?

OpenStudy (anonymous):

Haha youre not. its all about practice :)

OpenStudy (anonymous):

Wow john youre working hard :) Is that a question aimed at us or bubbles?:P

OpenStudy (johnweldon1993):

Right everyone learns at a different pace and sometimes it takes more than 1 way of looking at a problem to get it :)

OpenStudy (johnweldon1993):

Lol ehh in general ...everyone! lol :D

OpenStudy (anonymous):

now you distribute to simplify ? :o

OpenStudy (anonymous):

oh. then you equate the sides to get a quadratic equation, use the discriminant formula, get the x, take the positive value, and add 15 to it :P and there you have it, the asnwer :P

OpenStudy (johnweldon1993):

Mmhmm...distribute everything you can...what do you get @bubbles-are-cool.

OpenStudy (anonymous):

4x+60+4x = x^2 +15x and if when you add the 4x you get : 8x+60 = x^2 +15x

OpenStudy (johnweldon1993):

Right...so now you have that \[\large 8x + 60 = x^2 + 15x\] Now any terms you can combine?

OpenStudy (anonymous):

8x and 15x?

OpenStudy (johnweldon1993):

Mmhmm you can subtract 8x from both sides... \[\large x^2 + 7x = 60\] You can see this is turning into a quadratic...so go ahead and subtract 60 from both sides too \[\large x^2 + 7x - 60 = 0\]

OpenStudy (johnweldon1993):

Now you just use the quadratic equation to solve for your 'x' :)

OpenStudy (anonymous):

How would i do that if i have an exponent?

OpenStudy (johnweldon1993):

Well remember the quadratic equation uses the coefficients of this equation.. in the form \[\large ax^2 + bx + c = 0\] The coefficients are 'a' 'b' and 'c' right? So here... \[\large x^2 + 7x - 60\] a = 1 b = 7 c = -60 Make sense?

OpenStudy (anonymous):

yes c:

OpenStudy (johnweldon1993):

Perfect :) So you plug them into the quadratic equation...just in case you need to see it again... \[\large \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\]

OpenStudy (anonymous):

yes & now i have -7 +- (sq289) / 2

OpenStudy (anonymous):

do i find the squareroot?

OpenStudy (johnweldon1993):

Yes you do :)

OpenStudy (anonymous):

now i have -7 +- 17 / 2 would i add or subtract next?

OpenStudy (johnweldon1993):

Well now...you have 2 equations... \[\large \frac{-7 + -17}{2}\] and \[\large \frac{-7 - (-17)}{2}\] Solve both, what do you get?

OpenStudy (anonymous):

the first answer is -12 and the second is 5

OpenStudy (johnweldon1993):

Perfect... Now we know...time isn't negative....so we can discard that -12 okay? So our x that we solved is 5....(not the answer yet) but ready for the last step?

OpenStudy (anonymous):

yes next i add 15?

OpenStudy (johnweldon1993):

Haha perfect beat me to it :)

OpenStudy (anonymous):

so the answer is 20 c:

OpenStudy (johnweldon1993):

That is correct :)

OpenStudy (anonymous):

thanks !

OpenStudy (johnweldon1993):

Anytime :) Great Work @bubbles-are-cool.